The December 31, 2019 balance sheet of Frost Company showed the following balances:
Truck                90,000
Accumulated depreciation: truck                38,000
On January 1, 2020, after a slight mishap, the company revises the estimates on this truck.
     Revised remaining useful life 2 years
     Revised salvage value                  1,000
INSTRUCTIONS: Compute the revised annual depreciation expense for this equipment.
On October 1, 2019, Beaufort Mining Inc. purchased a new mine.
Cost of the mine              640,000
Estimated tons of ore in this mine              500,000 tons
Ore mined and sold during 2017                80,000 tons
INSTRUCTIONS: Prepare the journal entry to record depletion expense for 2019
GENERAL JOURNAL
Date Account Title DEBIT CREDIT
Answer #1

Book value of Truck = Cost price – Accumulated depreciation

= 90,000 - 38,000

= $52,000

New depreciable cost = Book value of Truck – Revised residual value

= 52,000 - 1,000

= $51,000

Revised annual depreciation = New depreciable cost/Remaining useful life

= 51,000/2

= $25,500

Next question

Cost of the mine = $640,000

Estimated tons of ore in this mine = 500,000 tons

Ore mined and sold during 2019 = 80,000 tons

Depletion expense per ton = Cost of the mine/Estimated tons of ore

= 500,000/640,000

= $0.78125

Depletion expense for 2019 = Depletion expense per ton x Ore mined and sold during 2019

= 0.78125 x 80,000

= $62,500

Journal

Dec. 31, 2019 Depletion expense 62,500
Accumulated depletion - Ore mine 62,500

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.